# Physics-informed, boundary-constrained Gaussian process regression for the reconstruction of fluid flow fields

### v1

@@ Abstract
Gaussian process regression techniques have been used in fluid mechanics for the reconstruction of flow fields from a reduction-of-dimension perspective. A main ingredient in this setting is the construction of adapted covariance functions, or kernels, to obtain such estimates. In this paper, we derive physics-informed kernels for simulating two-dimensional velocity fields of an incompressible (divergence-free) flow around aerodynamic profiles. These kernels allow to define Gaussian process priors satisfying the incompressibility condition and the prescribed boundary conditions along the profile in a continuous manner. Such physical and boundary constraints can be applied to any pre-defined scalar kernel in the proposed methodology, which is very general and can be implemented with high flexibility for a broad range of engineering applications. Its relevance and performances are illustrated by numerical simulations of flows around a cylinder and a NACA 0412 airfoil profile, for which no observation at the boundary is needed at all.
@@ 1 Introduction
Complex physical phenomena such as fluid flows have been extensively studied from different perspectives, by using both deterministic and probabilistic approaches [10, 11, 13, 16, 23, 37]. The development of various techniques for measuring physical quantities in flows [46] such as velocity, vorticity, pressure, or Lagrangian particle motion [30, 31], as well as the availability of high-precision numerical simulations, offer a variety of information from different nature for the description of fluid dynamics. In the last decade, there has been a strong interest in developing techniques that can be able to model these dynamics based on information from different sources; see for example [40, 45]. Nevertheless, the classical deterministic methods based on mesh construction (such as finite differences, finite volumes, or finite elements) are not necessarily adaptable for considering multiple-source numerical estimates and assimilating the available information if any. In the present paper, we are interested in the reconstruction of physical fields from a two-dimensional (2D) incompressible flow around an aerodynamic profile. The constructed estimates must be geometrically flexible and predict the physical fields at any unobserved spatial position based on a set of discrete direct observations of the field. Moreover, these estimates have to simultaneously satisfy continuous-type conditions in agreement with the physics of the flow.
@@
To achieve this, we use a Gaussian process (GP) framework by defining a priori distributions of the physical fields, and imposing incompressibility and profile boundary conditions to them. This is done by modelling the scalar stream function of the fluid flow as a boundary-constrained GP, so that its velocity counterpart is a physics-informed prior that already fulfills the foregoing conditions in a continuous sense at all positions. The approach we use is based on a spectral projection of a prescribed GP onto a localized compact set, as derived in [17]. By this means, the physical conditions are imposed over a continuous set rather than discretely at specific observation points. Ultimately, these defined priors can be used in Gaussian process regression (GPR) based on pointwise observations of the fluid flow fields. This procedure is a relevant meshfree approach describing the fluid dynamics that can be directly merged within Lagrangian-based estimates [33]. Moreover, it is general enough since it is presented without restrictions on the geometry of the localized compact set that may correspond to the profile boundary in the applications. Additionally, an important aspect of using GP for the reconstruction of velocity and vorticity fields is the fact that we do not need to solve a pressure equation in order to get estimates of these fields.
@@
In order to completely define a GP, a mean function and a covariance function (or a kernel as defined in the dedicated literature) are required. These functions define the mathematical properties of the prior estimate, thus yielding an automatic satisfaction of the problem constraints if desired. The nature of these constraints can be set by physical specifications of the modelling problem, still no unique rule-of-thumb exists for incorporating them in the mean and/or kernel functions. Specific techniques to learn them from the data have been proposed in e.g. [1, 35]. Divergence-free random vector fields have been considered in [42] and used for Lagrangian flow particle simulation in [33]. Additionally, interests in constraining GP priors to boundary conditions on the boundary of domains with different geometries have also emerged [12, 19, 44]. In [19], the boundary-constrained GPs are built from a particular structure based on Fourier expansions, whereas [44] presents a spectral framework based on eigenfunctions of the Laplacian that is applicable to stationary base kernels. Other methods, such as [12], define boundary-constrained GP kernels from multiplication of the base kernel with a symmetric function satisfying the boundary conditions. We present a general method, developed firstly by [17], that modifies any given GP with a continuous kernel to satisfy a null condition (also known as homogeneous Dirichlet condition in the context of partial differential equations (PDEs)) over a compact subset of its domain.
@@
For a bounded subset $B\subset\mathbb{R}^{n}$, the $n$-dimensional volume of $B$ is denoted by $\left\lvert B\right\rvert$. $\mathcal{M}_{m,n}(\mathbb{R})$ stands for the set of $m\times n$ real matrices, with $\mathcal{M}_{n}(\mathbb{R})=\mathcal{M}_{n,n}(\mathbb{R})$. $\operatorname{Tr}{\boldsymbol{A}}$ stands for the trace of the square matrix ${\boldsymbol{A}}$. The usual Euclidean norm of a vector ${\boldsymbol{a}}$ in $\mathbb{R}^{n}$ is denoted by $\left\|{\boldsymbol{a}}\right\|$, and the scalar product of two vectors ${\boldsymbol{a}}$ and ${\boldsymbol{b}}$ in $\mathbb{R}^{n}$ is denoted by ${\boldsymbol{a}}\cdot{\boldsymbol{b}}$. Their tensor product is denoted by ${\boldsymbol{a}}\otimes{\boldsymbol{b}}$, such that $({\boldsymbol{a}}\otimes{\boldsymbol{b}}){\boldsymbol{c}}=({\boldsymbol{b}}\cdot{\boldsymbol{c}}){\boldsymbol{a}}$ for any ${\boldsymbol{c}}\in\mathbb{R}^{n}$. ${\boldsymbol{\nabla}}$ and curl stand for the usual gradient and curl vector operators, respectively, which are also denoted by $\smash{{\boldsymbol{\nabla}}_{\boldsymbol{x}}}$ and $\smash{{\textbf{curl}}_{\boldsymbol{x}}}$, respectively, to make explicit that they are computed with respect to the coordinates ${\boldsymbol{x}}$ if necessary. We denote by ${\boldsymbol{a}}^{\perp}$ the counterclockwise orthogonal vector to ${\boldsymbol{a}}\in\mathbb{R}^{2}$, such that ${\boldsymbol{a}}\cdot{\boldsymbol{a}}^{\perp}={\boldsymbol{a}}^{\perp}\cdot{\boldsymbol{a}}=0$. For ${\boldsymbol{x}}=(x_{1},x_{2})\in\mathbb{R}^{2}$ we have ${\textbf{curl}}_{\boldsymbol{x}}=\smash{{\boldsymbol{\nabla}}_{\boldsymbol{x}}^{\perp}}=\smash{(-\partial_{x_{2}},\partial_{x_{1}})^{\mathsf{T}}}$, where $\mathsf{T}$ stands for the transpose operator. For real-valued, second-order random variables ${Z}$ and ${Z}^{\prime}$ defined on a probability space $(X,{\mathcal{F}},P)$, their mathematical expectations, or averages, are denoted by $\mathbb{E}\left\{{Z}\right\}$ and $\mathbb{E}\left\{{Z}^{\prime}\right\}$, respectively, their covariance is denoted by $\operatorname{Cov}\left\{{Z},{Z}^{\prime}\right\}=\mathbb{E}\left\{({Z}-\mathbb{E}\left\{{Z}\right\})({Z}^{\prime}-\mathbb{E}\left\{{Z}^{\prime}\right\})\right\}$, and their variances are denoted by $\operatorname{Var}\left\{{Z}\right\}=\operatorname{Cov}\left\{{Z},{Z}\right\}$ and $\operatorname{Var}\left\{{Z}^{\prime}\right\}=\operatorname{Cov}\left\{{Z}^{\prime},{Z}^{\prime}\right\}$, respectively. A Gaussian process $\{{Z}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\}\sim{\mathcal{GP}}(m,{G})$ indexed on some set ${\Omega}$ is defined as the collection of Gaussian random variables ${Z}({\boldsymbol{x}})$ such that $\mathbb{E}\left\{{Z}({\boldsymbol{x}})\right\}=m({\boldsymbol{x}})$ and $\operatorname{Var}\left\{{Z}({\boldsymbol{x}}),{Z}({\boldsymbol{x}})\right\}={G}({\boldsymbol{x}},{\boldsymbol{x}})$, where ${\boldsymbol{x}}\mapsto m({\boldsymbol{x}})$ is their mean function and ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\mapsto{G}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})$ is their covariance function. Here we note ${Z}\sim{\mathscr{L}}$ if the random variable ${Z}$ is distributed according to the probability law ${\mathscr{L}}$. The classical Gaussian probability density function on $\mathbb{R}$ is $f_{{\mathcal{N}}}(z;m,\sigma^{2})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp[-\frac{1}{2}(\frac{z-m}{\sigma})^{2}]$.
@@ 2 Reminder on Gaussian Process Regression
We first recall the general setting of GPR from a probabilistic and functional approach. We refer to [3, 34, 36], for example, for a general introduction on reproducing kernel Hilbert spaces (RKHS) and kernel methods. Let us consider a domain ${\Omega}\subset\mathbb{R}^{d}$ where $d$ is a positive integer. Given a scalar-valued positive semi-definite symmetric function, or kernel ${G}:{\Omega}\times{\Omega}\to\mathbb{R}$, denote by ${{\mathcal{H}}_{G}}$ the RKHS associated to ${G}$, with induced inner product $\left\langle\cdot,\cdot\right\rangle_{G}$. By definition, we have the following properties:
@@
We will denote as ${Z}=\{{Z}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\}\sim{\mathcal{GP}}\left(0,{G}\right)$ the zero-mean Gaussian process indexed on ${\boldsymbol{x}}\in{\Omega}$, such that its covariance function is the kernel ${G}$. This function defines the structure of the sample paths of the GP in a certain extent, even though these sample paths do not necessarily belong to the RKHS associated to the kernel [24].
Now, consider a set of $N$ points ${\boldsymbol{X}}=\smash{\{{\boldsymbol{x}}_{i}\}_{i=1}^{N}}$ inside ${\Omega}$, and a vector ${\boldsymbol{Y}}=\smash{(y_{1},\ldots,y_{N})^{\mathsf{T}}}$ of $N$ fixed real-valued observations of ${Z}$ at these points. A generic interpolation formula to estimate the value of ${Z}({\boldsymbol{x}})$ at any unobserved ${\boldsymbol{x}}\in{\Omega}$ conditioned on the information that ${Z}({\boldsymbol{x}}_{i})=y_{i}$ for all $i=1,\ldots,N$ (denoted by ${Z}({\boldsymbol{X}})={\boldsymbol{Y}}$) is given by the representer theorem [34]:
@@
Here ${G}({\boldsymbol{X}},{\boldsymbol{X}})$ is the Gram matrix with elements $[{G}({\boldsymbol{X}},{\boldsymbol{X}})]_{ij}={G}({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j})$ and ${G}({\boldsymbol{x}},{\boldsymbol{X}})$ is the column vector with elements ${G}({\boldsymbol{x}},{\boldsymbol{X}})_{i}={G}({\boldsymbol{x}},{\boldsymbol{x}}_{i})$. Moreover, the counterpart for quantifying the covariance of the random variables ${Z}({\boldsymbol{x}})$ and ${Z}({\boldsymbol{x}}^{\prime})$ associated to the locations ${\boldsymbol{x}}$ and ${\boldsymbol{x}}^{\prime}$ in the domain ${\Omega}$, conditioned on ${Z}({\boldsymbol{X}})={\boldsymbol{Y}}$, reads:
@@
For a kernel ${G}$ that is regular enough, we can access to its pointwise derivatives and thus obtain estimates based on derivative information of the underlying target or phenomena. In this case, considering $M$ data points ${\boldsymbol{Q}}=\smash{\left\{{\boldsymbol{q}}_{i}\right\}_{i=1}^{M}}$ in ${\Omega}$ for the observation of derivatives, we can estimate ${Z}({\boldsymbol{x}})$ for any ${\boldsymbol{x}}\in{\Omega}$, conditioned on the information $\mathcal{L}_{i}{Z}({\boldsymbol{q}}_{i})=v_{i}$, $i=1,\dots,M$, for the fixed real values in the vector ${\boldsymbol{V}}=\smash{(v_{1},\dots,v_{M})^{\mathsf{T}}}$ and the corresponding linear differential operators $\mathcal{L}_{i}$:
[…]
Here $\smash{\mathcal{L}\otimes\mathcal{L}^{\prime}{G}({\boldsymbol{Q}},{\boldsymbol{Q}})}$ is the Gram matrix of entries $\mathcal{L}_{i}\otimes\mathcal{L}_{j}^{\prime}{G}({\boldsymbol{q}}_{i},{\boldsymbol{q}}_{j})$ for $i,j=1,\dots,M$, and $\smash{\mathcal{L}_{i}}$, $\smash{\mathcal{L}_{j}^{\prime}}$ refer to the differential operators acting on the first and second entries of ${G}$, respectively. The order of these differential operators is limited by the regularity of $G$. The representation identity of Eq. (2.3) remains valid if we replace the differential operators $\mathcal{L}_{i}$ with any other linear operators acting on the GP [18] since formally any linear evaluation of a GP remains a GP. The covariance counterpart for ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$ reads:
[…]
@@ 3.1 Constraints over a compact set
For a compact domain ${\Omega}\subset\mathbb{R}^{d}$ and arbitrary scalar GP ${Z}=\{{Z}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\}\sim{\mathcal{GP}}\left(0,{G}\right)$ with covariance continuous kernel ${G}:{\Omega}\times{\Omega}\rightarrow\mathbb{R}$, we now detail how to modify ${G}$ such that the resulting zero-mean GP satisfies homogeneous Dirichlet boundary conditions on a compact set $\Gamma$ inside ${\Omega}$. For a non-homogeneous condition, the resulting constrained kernel can be associated to a GP with a mean function $m$ satisfying the desired Dirichlet boundary condition. The results presented in this section are based on [17].
Firstly, recall the Karhunen-Loève expansion of ${Z}$ with respect to a particular measure ${\nu}$ over the compact set $\Gamma$:
@@
The measure ${\nu}$ can be chosen freely as any probability measure over $\Gamma$. The eigenvalues of $\mathcal{T}_{\Gamma}$ are then real, positive, and can be ordered in a decreasing sequence converging to 00.
@@
Consider an arbitrary scalar Gaussian process ${Z}=\{{Z}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\}\sim{\mathcal{GP}}\left(0,{G}\right)$, with continuous covariance kernel ${G}:{\Omega}\times{\Omega}\rightarrow\mathbb{R}$, and a compact subset $\Gamma$ of its domain ${\Omega}$. Let $\lambda_{n},{\phi}_{n}$ and $\xi_{n}$, $n\in\mathbb{N}$, come from the Karhunen-Loève expansion of ${Z}$ on $\Gamma$ with respect to a measure ${\nu}$, and $\tilde{{\phi}}_{n}$ be the lifting of ${\phi}_{n}$ to ${\Omega}$ as defined in Eq. (3.3). Then ${{Z}_{0}}=\left\{{{Z}_{0}}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\right\}$ defined as:
@@
This is immediately verified since, if ${\boldsymbol{x}}\in\Gamma$, the lifting $\tilde{{\phi}}_{n}$ coincides with ${\phi}_{n}$, implying that the term $\sum_{n=0}^{+\infty}\sqrt{\lambda_{n}}\,\tilde{{\phi}}_{n}({\boldsymbol{x}})\,\xi_{n}$ coincides with the Karhunen-Loève expansion of ${Z}$ at ${\boldsymbol{x}}$.

### v2

@@ Abstract
Gaussian process regression techniques have been used in fluid mechanics for the reconstruction of flow fields from a reduction-of-dimension perspective. A main ingredient in this setting is the construction of adapted covariance functions, or kernels, to obtain such estimates. In this paper, we present a general method for constraining a prescribed Gaussian process on an arbitrary compact set. The kernel of the pre-defined process must be at least continuous and may include other information about the studied phenomenon. This general boundary-constraining framework can be implemented with high flexibility for a broad range of engineering applications. From this, we derive physics-informed kernels for simulating two-dimensional velocity fields of an incompressible (divergence-free) flow around aerodynamic profiles. These kernels allow to define Gaussian process priors satisfying the incompressibility condition and the prescribed boundary conditions along the profile in a continuous manner. We describe an adapted numerical method for the boundary-constraining procedure parameterized by a measure on the compact set. The relevance of the methodology and performances are illustrated by numerical simulations of flows around a cylinder and a NACA 0412 airfoil profile, for which no observation at the boundary is needed at all.
@@ 1 Introduction
Complex physical phenomena such as fluid flows have been extensively studied from different perspectives, by using both deterministic and probabilistic approaches [10, 11, 13, 16, 23, 38]. The development of various techniques for measuring physical quantities in flows [47] such as velocity, vorticity, pressure, or Lagrangian particle motion [30, 31], as well as the availability of high-precision numerical simulations, offer a variety of information from different nature for the description of fluid dynamics. In the last decade, there has been a strong interest in developing techniques that can be able to model these dynamics based on information from different sources; see for example [41, 46]. Nevertheless, the classical deterministic methods based on mesh construction (such as finite differences, finite volumes, or finite elements) are not necessarily adaptable for considering multiple-source numerical estimates and assimilating the available information if any. In the present paper, we are interested in developing a physics-informed Gaussian process (GP) approach for the reconstruction of physical fields of an incompressible fluid flow around an aerodynamic profile. We present a general methodology for constraining a prescribed GP on an arbitrary compact set, which corresponds to the boundary of the profile in the applications. The pre-defined covariance function may include other physical information and is only required to be continuous on the compact set for performing the constraining procedure. In particular, we use this framework for the reconstruction of two-dimensional (2D) incompressible flows around a cylinder and a NACA aerodynamic profile. The obtained estimates must predict the physical fields at any unobserved spatial position based on a set of discrete direct observations of the field. Moreover, they have to simultaneously satisfy continuous-type conditions in agreement with the physics of the flow.
@@
To achieve this, we firstly describe in a general context the spectral GP constraining method (initially developed in [17]). This presentation is intended such that the proposed approach can be used in other engineering applications involving boundary constraints under general geometries. We also describe a numerical method for solving the associated eigenvalue problem, which is parameterized by a probability measure defined on the compact set. We study two choices for the definition of this measure. Next, we define our GP modelling framework by imposing incompressibility and profile boundary conditions to a priori distributions of the physical fields. This is done by modelling the scalar stream function of the fluid flow as a boundary-constrained GP, so that its velocity counterpart is a physics-informed prior that already fulfills the foregoing conditions in a continuous sense at all positions. Namely, these constraints are imposed over a continuous set rather than discretely at specific observation points. Ultimately, these defined priors can be used in Gaussian process regression (GPR) based on pointwise observations of the fluid flow fields. The main advantage of this approach is that there is no need to save some part of the budget of the design of experiments for the inclusion of the physical constraints as observations since they are already included in the a priori distributions. Furthermore, this procedure is a relevant mesh-free approach describing the fluid dynamics that can be directly merged within Lagrangian-based estimates [34]. Additionally, this framework can prevent wrong predictions of Lagrangian particles moving inside the aerodynamic profile, which may occur when using discrete observations of the boundary condition instead. Moreover, an important aspect of using GPs for the reconstruction of velocity and vorticity fields is the fact that we do not need to solve a pressure equation in order to get estimates of these fields.
@@
In order to completely define a GP, a mean function and a covariance function (or a kernel as defined in the dedicated literature) are required. These functions define the mathematical properties of the prior estimate, thus yielding an automatic satisfaction of the problem constraints if desired. The nature of these constraints can be set by physical specifications of the modelling problem, still no unique rule-of-thumb exists for incorporating them in the mean and/or kernel functions. Specific techniques to learn them from the data have been proposed in e.g. [1, 36]. Divergence-free random vector fields have been considered in [43] and used for Lagrangian flow particle simulation in [34]. Additionally, interests in constraining GP priors to boundary conditions on the boundary of domains with different geometries have also emerged [12, 19, 45]. In [19], the boundary-constrained GPs are built from a particular structure based on Fourier expansions, whereas [45] presents a spectral framework based on eigenfunctions of the Laplacian that is applicable to stationary base kernels. Other methods, such as [12], define boundary-constrained GP kernels from multiplication of the base kernel with a symmetric function satisfying the boundary conditions. Our chosen method, developed firstly by [17], modifies any given GP with a continuous kernel to satisfy a null condition (also known as homogeneous Dirichlet condition in the context of partial differential equations (PDEs)) over any compact subset of its domain. We provide an associated numerical method for scalability of this procedure under different choices of the probability measure. To our knowledge, this is the first time that such a general boundary-constraining method (adapted for computing kernel derivatives) is applied to an arbitrary continuous kernel (not only stationary) and under a complex definition of the arbitrary constraining set, such as the boundary of a NACA airfoil.
@@
[…]
@@ 2 Reminder on Gaussian Process Regression
We first recall the general setting of GPR from a probabilistic and functional approach. We refer to [3, 35, 37], for example, for a general introduction on reproducing kernel Hilbert spaces (RKHS) and kernel methods. Let us consider a domain ${\Omega}\subset\mathbb{R}^{d}$ where $d$ is a positive integer. Given a scalar-valued positive semi-definite symmetric function, or kernel ${G}:{\Omega}\times{\Omega}\to\mathbb{R}$, denote by ${\mathcal{H}}_{G}$} the RKHS associated to G{G}, with induced inner product $\left\langle\cdot,\cdot\right\rangle_{G}$. By definition, we have the following properties:
@@
We will denote as ${Z}=\{{Z}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\}\sim{\mathcal{GP}}\left(0,{G}\right)$ the zero-mean Gaussian process indexed on ${\boldsymbol{x}}\in{\Omega}$, such that its covariance function is the kernel G{G}. This function defines the structure of the sample paths of the GP in a certain extent, even though these sample paths do not necessarily belong to the RKHS associated to the kernel [24].
Now, consider a set of $N$ points ${\boldsymbol{X}}=\smash{\{{\boldsymbol{x}}_{i}\}_{i=1}^{N}}$ inside $\Omega$}, and a vector ${\boldsymbol{Y}}=\smash{(y_{1},\ldots,y_{N})^{\mathsf{T}}}$ of $N$ fixed real-valued observations of Z{Z} at these points. A generic interpolation formula to estimate the value of ${Z}({\boldsymbol{x}})$ at any unobserved ${\boldsymbol{x}}\in{\Omega}$ conditioned on the information that ${Z}({\boldsymbol{x}}_{i})=y_{i}$ for all $i=1,\ldots,N$ (denoted by ${Z}({\boldsymbol{X}})={\boldsymbol{Y}}$) is given by the representer theorem [35]:
@@
Here ${G}({\boldsymbol{X}},{\boldsymbol{X}})$ is the Gram matrix with elements ${G}({\boldsymbol{X}},{\boldsymbol{X}})]_{ij}={G}({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j})$ and ${G}({\boldsymbol{x}},{\boldsymbol{X}})$ is the column vector with elements ${G}({\boldsymbol{x}},{\boldsymbol{X}})_{i}={G}({\boldsymbol{x}},{\boldsymbol{x}}_{i})$. Moreover, the counterpart for quantifying the covariance of the random variables ${Z}({\boldsymbol{x}})$ and ${Z}({\boldsymbol{x}}^{\prime})$ associated to the locations $\boldsymbol{x}$} and $\boldsymbol{x}$$^{\prime}$ in the domain $\Omega$}, conditioned on ${Z}({\boldsymbol{X}})={\boldsymbol{Y}}$, reads:
@@
For a kernel G{G} that is regular enough, we can access to its pointwise derivatives and thus obtain estimates based on derivative information of the underlying target or phenomena. In this case, considering $M$ data points ${\boldsymbol{Q}}=\smash{\left\{{\boldsymbol{q}}_{i}\right\}_{i=1}^{M}}$ in $\Omega$} for the observation of derivatives, we can estimate ${Z}({\boldsymbol{x}})$ for any ${\boldsymbol{x}}\in{\Omega}$, conditioned on the information $\mathcal{L}_{i}{Z}({\boldsymbol{q}}_{i})=v_{i}$, $i=1,\dots,M$, for the fixed real values in the vector ${\boldsymbol{V}}=\smash{(v_{1},\dots,v_{M})^{\mathsf{T}}}$ and the corresponding linear differential operators $\mathcal{L}_{i}$:
$\mathbb{E}\left\{{Z}({\boldsymbol{x}})\ |\ \mathcal{L}_{i}{Z}({\boldsymbol{q}}_{i})=v_{i},\ 1\leq i\leq M\right\}={\mathcal{L}^{\prime}}{G}({\boldsymbol{x}},{\boldsymbol{Q}})\left(\mathcal{L}\otimes\mathcal{L}^{\prime}{G}({\boldsymbol{Q}},{\boldsymbol{Q}})\right)^{-1}{\boldsymbol{V}}\,.$ (2.3)
Here $\smash{\mathcal{L}\otimes\mathcal{L}^{\prime}{G}({\boldsymbol{Q}},{\boldsymbol{Q}})}$ is the Gram matrix of entries $\mathcal{L}_{i}\otimes\mathcal{L}_{j}^{\prime}{G}({\boldsymbol{q}}_{i},{\boldsymbol{q}}_{j})$ for $i,j=1,\dots,M$, and $\smash{\mathcal{L}_{i}}$, $\smash{\mathcal{L}_{j}^{\prime}}$ refer to the differential operators acting on the first and second entries of G{G}, respectively. The order of these differential operators is limited by the regularity of $G$. The representation identity of Eq. (2.3) remains valid if we replace the differential operators $\mathcal{L}_{i}$ with any other linear operators acting on the GP [18] since formally any linear evaluation of a GP remains a GP. The covariance counterpart for ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$ reads:
$\operatorname{Cov}\left\{{Z}({\boldsymbol{x}}),{Z}({\boldsymbol{x}}^{\prime})\ |\ \mathcal{L}_{i}{Z}({\boldsymbol{q}}_{i})=v_{i},\ 1\leq i\leq M\right\}={G}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})-\mathcal{L}^{\prime}{G}({\boldsymbol{x}},{\boldsymbol{Q}})\left(\mathcal{L}\otimes\mathcal{L}^{\prime}{G}({\boldsymbol{Q}},{\boldsymbol{Q}})\right)^{-1}\mathcal{L}{G}({\boldsymbol{Q}},{\boldsymbol{x}}^{\prime})\,.$ (2.4)
@@ 3.1 Constraints over a compact set
For a compact domain ${\Omega}\subset\mathbb{R}^{d}$ and arbitrary scalar GP ${Z}=\{{Z}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\}\sim{\mathcal{GP}}\left(0,{G}\right)$ with covariance continuous kernel ${G}:{\Omega}\times{\Omega}\rightarrow\mathbb{R}$, we now detail how to modify G{G} such that the resulting zero-mean GP satisfies homogeneous Dirichlet boundary conditions on a compact set $\Gamma$ inside $\Omega$}. For a non-homogeneous condition, the resulting constrained kernel can be associated to a GP with a mean function $m$ satisfying the desired Dirichlet boundary condition. The results presented in this section are based on [17].
Firstly, recall the Karhunen-Loève expansion of Z{Z} with respect to a particular measure $\nu$} over the compact set $\Gamma$:
@@
The measure $\nu$} can be chosen freely as any probability measure over $\Gamma$. The eigenvalues of $\mathcal{T}_{\Gamma}$ are then real, positive, and can be ordered in a decreasing sequence converging to 0.
@@
Consider an arbitrary scalar Gaussian process ${Z}=\{{Z}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\}\sim{\mathcal{GP}}\left(0,{G}\right)$, with continuous covariance kernel ${G}:{\Omega}\times{\Omega}\rightarrow\mathbb{R}$, and a compact subset $\Gamma$ of its domain $\Omega$}. Let $\lambda_{n},{\phi}_{n}$ and $\xi_{n}$, $n\in\mathbb{N}$, come from the Karhunen-Loève expansion of Z{Z} on $\Gamma$ with respect to a measure $\nu$}, and $\tilde{{\phi}}_{n}$ be the lifting of ${\phi}_{n}$ to $\Omega$} as defined in Eq. (3.3). Then ${{Z}_{0}}=\left\{{{Z}_{0}}({\boldsymbol{x}}),{\boldsymbol{x}}\in{\Omega}\right\}$ defined as:
@@
This is immediately verified since, if ${\boldsymbol{x}}\in\Gamma$, the lifting $\tilde{{\phi}}_{n}$ coincides with ${\phi}_{n}$, implying that the term $\sum_{n=0}^{+\infty}\sqrt{\lambda_{n}}\,\tilde{{\phi}}_{n}({\boldsymbol{x}})\,\xi_{n}$ coincides with the Karhunen-Loève expansion of Z{Z} at $\boldsymbol{x}$}.